OpenStudy (anonymous):

What are the solutions of 4x2 + x = –3?

OpenStudy (anonymous):

Get all the terms on one side. So, \[4x^2 + x + 3 =0\] Then you need to factorise. You're either going to have (x + a)(4x + b) or (2x+c)(2x+d) You need to find a combination where ab or cd = 3 and where 4ax+bx or 2cx+2dx = x

OpenStudy (anonymous):

Seeing as the factors of 3 are only 1 and 3, a and b or c and d must be 1 and 3.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i get that so far

OpenStudy (anonymous):

If you put that into the second equation you get (2x+3)(2x+1)= 4x^2 + 8x + 3 so it's not that. So its either (4x+1)(x+3) or (4x+3)(x+1)

OpenStudy (anonymous):

i think it's (4x + 1)(x + 3)

OpenStudy (anonymous):

Wait, they don't work either, are you sure it's not 4x^2 + x = 3?

OpenStudy (anonymous):

it must be then

OpenStudy (anonymous):

i know that the answer is going to be an imaginary number and a radical all over an integer

OpenStudy (anonymous):

i got -1 but i can be wrong

OpenStudy (anonymous):

Right well put it into the quadratic formula.\[x =(- b +- \sqrt{b^2-4ac} )/2a\]

OpenStudy (anonymous):

b = 1, a = 4 and c = 3.

OpenStudy (anonymous):

im working it out on paper to see if i can get it

OpenStudy (anonymous):

Ok :P This equation doesn't have any real roots so they'll both be imaginary.

OpenStudy (anonymous):

I got -1 + or - i√47 all over 8

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

No. You can't have 1 real root and 1 imaginary. Either 2 real or 2 imaginary.

OpenStudy (anonymous):

Oh right!! Yeah that's right.

OpenStudy (anonymous):

The 'or' confused me.

OpenStudy (anonymous):

wait so whats the final answer?

OpenStudy (anonymous):

\[\frac{ -1 + \sqrt{47}i }{ 8 }\]

OpenStudy (anonymous):

Or\[\frac{ -1 - \sqrt{47}i }{ 8 }\]

OpenStudy (anonymous):

i being root -1 and therefore imaginary.

OpenStudy (anonymous):

kk